OpenStudy (anonymous):

Find the midpoint of the line segment whose endpoints are (5, 9) and (-3, 7).

OpenStudy (ivettef365):

use midpoint formula

jimthompson5910 (jim_thompson5910):

to find the x coordinate of the midpoint, you add up the two x coordinates and then divide by 2

OpenStudy (anonymous):

which is

OpenStudy (ivettef365):

http://cs.selu.edu/~rbyrd/math/midpoint/

jimthompson5910 (jim_thompson5910):

so xmid = (x1 + x2)/2 xmid = (5 + (-3))/2 xmid = 2/2 xmid = 1 which means the x coordinate of the midpoint is x = 1

jimthompson5910 (jim_thompson5910):

do the same for the y coordinates

OpenStudy (anonymous):

thanks

OpenStudy (ivettef365):

did you try the website to find midpoint ?

OpenStudy (anonymous):

yes

OpenStudy (ivettef365):

:)

OpenStudy (anonymous):

Determine the solution set of ax 2 + bx + c = 0.